If h(t)=t^2-2t find the following:
h(2)
H(-1)
h(x+2)

Answer by checkley77(12844)

h(t)=t^2-2t
h(2)
H(2)=2^2-2*2
H(2)=4-4=0 ANS.
----------------
H(-1)
H(-1)=1^2-2*1
H(-1)=1-2=-1 ANS.
-------------------
H(x+2)
H(x+2)=(x+2)^2-2(x+2)
H(x+2)=x^2+4x+4-2x-4
H(x+2=x^2+2x
H(x+2)=x(x+2)